# Astra[Uncaptioned image]: General Interactive World Model with Autoregressive Denoising

### v2

@@ 4.4 Extended applications
Autonomous driving. We extend Astra to autonomous driving using the nuScenes dataset (caesar2020nuscenes), which provides multi-view videos and diverse traffic contexts. Given ego-vehicle observations and discrete control actions (e.g., turn left, move forward), Astra generates realistic future driving videos that capture vehicle motion, road geometry, and agent interactions. Its combination of long-term coherence and precise action responsiveness enables interactive driving simulation.
@@ A.1 Datasets
We use the training data sampled from the nuScenes (caesar2020nuscenes), Sekai (li2025sekai), SpatialVID (wang2025spatialvid), Multi-Cam Video (bai2025recammaster) and RT-1 (brohan2022rt). To ensure balanced exposure across these datasets, we apply a set of sampling weights that control the frequency with which data from each source is drawn during training. A summary of dataset statistics is provided in Table 1, with detailed descriptions reported as follows:
@@
RT-1: RT-1 is a large, diverse dataset of robot trajectories that includes multiple tasks, objects and environments. It contains over 130k individual demonstrations constituting over 700 distinct task instructions using a large variety of objects. We use its 7 dimensions ee-space pose parts as our action. We use Open X-Embodiment (o2024open) version of the RT-1 dataset.
@@ A.2 Training details
We initialize our model using the pre-trained weights of the Wan-2.1 base model (wan2025wan). Our training is run on 8 GPUs (80G) with a per-GPU batch size of 1. We train our model with AdamW (loshchilov2017decoupled) optimizer and a learning rate of 1e51e{-5} for 30 epochs. Our training runs take approximately 24 hours to converge. All videos are resized and cropped to the training resolution (480 × 832). The model is trained on the latent space produced by a 3D VAE. In pixel space, the count of condition frames is randomly sampled from the range [1, 128], whereas the number of target frames is consistently set to 33.
@@ A.3 Model architecture
Our framework builds on Wan-2.1 (wan2025wan), a large-scale video diffusion model composed of 30 stacked flow transformer (DiT-style) blocks, each containing multi-head self-attention and feed-forward layers with residual connections. Wan-2.1 serves as a strong pre-trained backbone, providing rich generative priors for high-quality video synthesis. To enable interactive control, we introduce two lightweight yet effective extensions: ACT-Adapter and Mixture of Action Experts (MoAE). The ACT-Adapter is implemented as a single linear layer inserted after each self-attention block. It is initialized as an identity mapping and fine-tuned jointly with the attention parameters, allowing the model to inject action features into the latent space while preserving stability of the pre-trained weights. In parallel, MoAE provides a modular action encoding mechanism. It consists of a linear router that projects heterogeneous action modalities into a shared space, followed by a set of MLP-based experts specialized for different action types. Specifically, we support camera control actions represented as 7 or 12-dimensional vectors, robotic actions represented as 7-dimensional vectors, and navigation commands expressed through discrete keyboard and mouse inputs. The routed outputs are combined into a unified representation that is fed into the ACT-Adapter, ensuring both precision and versatility. Together, these components augment Wan-2.1 into an autoregressive denoising model capable of faithfully following actions while maintaining long-horizon temporal coherence. During inference, we set the scale $s$ for action-free guidance (AFG) to be 3.0.
@@
Wan-2.1 (wan2025wan) 2.96 7.37 0.061 0.691
YUME (mao2025yume) 2.20 5.80 0.268 0.741
MatrixGame (he2025matrix) 2.25 5.63 0.652 0.748
NWM (bar2025navigation) 2.47 6.13 0.311 0.635
@@ A.4 Metric details
We follow YUME (mao2025yume) to use human evaluation and VBench (huang2024vbench) metrics to assess the performance of our model. For instruction following, automated camera-motion estimators (e.g., MegaSaM (li2025megasam)) can provide approximate motion labels, but their predictions often suffer from inaccuracies and quantization errors, especially under complex scene geometry or fast motion. To ensure reliable assessment, we therefore rely purely on human evaluation. Specifically, we recruit 20 users to inspect each generated sequence together with the corresponding action command. The instruction-following score is computed as the ratio of users who agree that the generated motion faithfully reflects the specified action direction and action type. For a more objective evaluation, we follow prior works (bai2025recammaster; he2024cameractrl) and measure how closely the camera motion in the generated video matches the ground-truth trajectory, using MegaSaM to estimate camera poses. As shown in Table A, Astra achieves lower rotation and translation errors than existing methods, consistent with the human-rated instruction-following results. Together, these metrics offer a fine-grained quantitative view of action alignment, further demonstrating the precise interactive control enabled by our model.
@@ A.5 Parameter analysis
Astra is designed to be lightweight, adding far fewer parameters than prior interactive world models such as YUME (mao2025yume) and MatrixGame (he2025matrix). As shown in Table B, MatrixGame introduces heavy cross-attention modules and large action encoders, resulting in significant training and inference overhead, while YUME depends on a much larger 13B backbone. In contrast, Astra adds only two small components: ACT-Adapter (a single linear layer after each self-attention block) and MoAE (a lightweight linear router plus small MLP experts, with only one expert active per step).
@@ Appendix B Comparison with existing world models
Table C compares four representative world models—Wan-2.1 (wan2025wan), MatrixGame (he2025matrix), YUME (mao2025yume), and Astra—across three key dimensions: domain, control modality, and interaction horizon. In terms of domain, Astra is designed for general-purpose scenarios, showcasing their versatility in a wide range of applications. MatrixGame, however, is tailored specifically for game-related contexts, limiting its use to game-specific environments. YUME focuses on walking-specific domains, which is a niche area compared to the general-purpose methods. For control modalities, Wan-2.1 relies solely on text input, reflecting a language-driven paradigm. MatrixGame and YUME adopt traditional keyboard and mouse controls. Astra, enabled by MoAE, supports multiple modalities—including camera input, keyboard/mouse, and robot pose—providing more flexible and intuitive interaction. Finally, regarding the interaction horizon, Wan-2.1 and MatrixGame are limited to short spans of a few seconds. YUME extends this to 8–10 seconds, and Astra matches this longer horizon through noisy memory and the input-packing technique of (zhang2025packinginputframecontext). Combined with its multi-modal control and general-purpose domain, Astra emerges as the most comprehensive solution among the compared methods.
@@
NVM (bar2025navigation) CDiT-XL $\sim$ 1B Full tuning
YUME (mao2025yume) Wan2.1-14B $\sim$ 14B Full tuning
MatrixGame (he2025matrix) Wan2.1-1.3B $\sim$ 1.8B Full tuning, cross-attn adapters
@@
Wan-2.1 (wan2025wan) General Text A few seconds
MatrixGame (he2025matrix) Game-specific
Keyboard / Mouse
A few seconds
YUME (mao2025yume) Walking-specific
Keyboard / Mouse
8-10 seconds
@@ C.2 Out-of-domain generalization
Figure B: Qualitative comparison with NVM (bar2025navigation). Astra consistently produces visually coherent results while faithfully following action inputs.
@@ C.3 Comparisons with more methods
To more comprehensively situate Astra within the broader landscape of visual world modeling, we provide extended quantitative and qualitative comparisons with additional state-of-the-art approaches. Specifically, we compare against methods such as Navigation World Models (NVM) (bar2025navigation) in Figure B and Table A, which are representative of recent approaches integrating multimodal inputs and action-conditioned predictions. Our results demonstrate that Astra achieves competitive or superior performance across key metrics, including visual fidelity, temporal consistency, and action responsiveness. While direct numerical comparisons are limited by differences in evaluation protocols and datasets, qualitative visualizations show that Astra generates long-horizon, coherent videos that faithfully follow user-specified actions, outperforming prior methods in capturing the causal dynamics of the environment. These comparisons highlight Astra’s ability to combine accurate action conditioning with long-term rollout stability, confirming its effectiveness as a general interactive world model across diverse scenarios. It is worth noting that many of the most advanced world models, such as Genie-3 (bruce2024genie), are currently closed-source or provide only limited evaluation interfaces, making direct numerical comparison infeasible.
@@ C.4 Quantitative comparisons on larger datasets
To further address concerns about evaluation scale, we conduct an expanded study on a larger held-out set from the CityWalker (liu2025citywalker) dataset. CityWalker is an egocentric urban navigation corpus collected from in-the-wild YouTube walking videos with pose estimates. We sample 100 scene images, each paired with its future ground-truth action trajectory, producing 100 long-horizon rollouts that robustly test cross-scene generalization and action adherence. Across this larger evaluation, Astra consistently achieves the best action-following accuracy and video quality, outperforming all baselines in metrics such as motion consistency, temporal stability, and VBench perceptual scores (Table D). These results confirm that Astra’s strong performance is not an artifact of small-set evaluation, and that the model maintains robust generalization and reliable action responsiveness even under substantially expanded test conditions.
@@
Method Instruction Following $\uparrow$ Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Wan-2.1 (wan2025wan) 0.084 0.827 0.843 0.913 0.417 0.632 MatrixGame (he2025matrix) 0.247 0.923 0.939 0.946 0.426 0.653 Yume (mao2025yume) 0.619 0.933 0.927 0.972 0.511 0.628 Astra (Ours) 0.641 0.948 0.944 0.983 0.554 0.695
@@ Appendix D Discussion on visual inertia
Previous work (zhang2025packinginputframecontext) has shown that longer video context can improve generation quality. However, in interactive world modeling, we observe that increasing context length can reduce action responsiveness. We term this phenomenon visual inertia—the model’s tendency to over-rely on past visual frames while neglecting user actions. To examine this, we train Astra with different history lengths without our proposed noisy-memory mechanism. As shown in Figure C, the action-following score drops substantially as context length increases, indicating that overly clean and long visual histories can dominate the model’s decision process. This motivates our noise-augmented memory design, which intentionally reduces visual dominance and encourages the model to integrate both historical context and action signals when generating future frames.

### v3

@@ 4.4 Extended applications
Autonomous driving. We extend Astra to autonomous driving using the nuScenes dataset (Caesar et al., 2020), which provides multi-view videos and diverse traffic contexts. Given ego-vehicle observations and discrete control actions (e.g., turn left, move forward), Astra generates realistic future driving videos that capture vehicle motion, road geometry, and agent interactions. Its combination of long-term coherence and precise action responsiveness enables interactive driving simulation.
@@ A.1 Datasets
We use the training data sampled from the nuScenes (Caesar et al., 2020), Sekai (Li et al., 2025a), SpatialVID (Wang et al., 2025), Multi-Cam Video (Bai et al., 2025) and RT-1 (Brohan et al., 2022). To ensure balanced exposure across these datasets, we apply a set of sampling weights that control the frequency with which data from each source is drawn during training. A summary of dataset statistics is provided in Table 1, with detailed descriptions reported as follows:
@@
RT-1: RT-1 is a large, diverse dataset of robot trajectories that includes multiple tasks, objects and environments. It contains over 130k individual demonstrations constituting over 700 distinct task instructions using a large variety of objects. We use its 7 dimensions ee-space pose parts as our action. We use Open X-Embodiment (O’Neill et al., 2024) version of the RT-1 dataset.
@@ A.2 Training details
We initialize our model using the pre-trained weights of the Wan-2.1 base model (Wan et al., 2025). Our training is run on 8 GPUs (80G) with a per-GPU batch size of 1. We train our model with AdamW (Loshchilov and Hutter, 2017) optimizer and a learning rate of 1e51e{-5} for 30 epochs. Our training runs take approximately 24 hours to converge. All videos are resized and cropped to the training resolution (480 × 832). The model is trained on the latent space produced by a 3D VAE. In pixel space, the count of condition frames is randomly sampled from the range [1, 128], whereas the number of target frames is consistently set to 33.
@@ A.3 Model architecture
Our framework builds on Wan-2.1 (Wan et al., 2025), a large-scale video diffusion model composed of 30 stacked flow transformer (DiT-style) blocks, each containing multi-head self-attention and feed-forward layers with residual connections. Wan-2.1 serves as a strong pre-trained backbone, providing rich generative priors for high-quality video synthesis. To enable interactive control, we introduce two lightweight yet effective extensions: ACT-Adapter and Mixture of Action Experts (MoAE). The ACT-Adapter is implemented as a single linear layer inserted after each self-attention block. It is initialized as an identity mapping and fine-tuned jointly with the attention parameters, allowing the model to inject action features into the latent space while preserving stability of the pre-trained weights. In parallel, MoAE provides a modular action encoding mechanism. It consists of a linear router that projects heterogeneous action modalities into a shared space, followed by a set of MLP-based experts specialized for different action types. Specifically, we support camera control actions represented as 7 or 12-dimensional vectors, robotic actions represented as 7-dimensional vectors, and navigation commands expressed through discrete keyboard and mouse inputs. The routed outputs are combined into a unified representation that is fed into the ACT-Adapter, ensuring both precision and versatility. Together, these components augment Wan-2.1 into an autoregressive denoising model capable of faithfully following actions while maintaining long-horizon temporal coherence. During inference, we set the scale $s$ for action-free guidance (AFG) to be 3.0.
@@
Wan-2.1 (Wan et al., 2025) 2.96 7.37 0.061 0.691
YUME (Mao et al., 2025) 2.20 5.80 0.268 0.741
MatrixGame (He et al., 2025) 2.25 5.63 0.652 0.748
NWM (Bar et al., 2025) 2.47 6.13 0.311 0.635
@@ A.4 Metric details
We follow YUME (Mao et al., 2025) to use human evaluation and VBench (Huang et al., 2024) metrics to assess the performance of our model. For instruction following, automated camera-motion estimators (e.g., MegaSaM (Li et al., 2025b)) can provide approximate motion labels, but their predictions often suffer from inaccuracies and quantization errors, especially under complex scene geometry or fast motion. To ensure reliable assessment, we therefore rely purely on human evaluation. Specifically, we recruit 20 users to inspect each generated sequence together with the corresponding action command. The instruction-following score is computed as the ratio of users who agree that the generated motion faithfully reflects the specified action direction and action type. For a more objective evaluation, we follow prior works (Bai et al., 2025; He et al., 2024) and measure how closely the camera motion in the generated video matches the ground-truth trajectory, using MegaSaM to estimate camera poses. As shown in Table A, Astra achieves lower rotation and translation errors than existing methods, consistent with the human-rated instruction-following results. Together, these metrics offer a fine-grained quantitative view of action alignment, further demonstrating the precise interactive control enabled by our model.
@@ A.5 Parameter analysis
Astra is designed to be lightweight, adding far fewer parameters than prior interactive world models such as YUME (Mao et al., 2025) and MatrixGame (He et al., 2025). As shown in Table B, MatrixGame introduces heavy cross-attention modules and large action encoders, resulting in significant training and inference overhead, while YUME depends on a much larger 13B backbone. In contrast, Astra adds only two small components: ACT-Adapter (a single linear layer after each self-attention block) and MoAE (a lightweight linear router plus small MLP experts, with only one expert active per step).
@@ Appendix B Comparison with existing world models
Table C compares four representative world models—Wan-2.1 (Wan et al., 2025), MatrixGame (He et al., 2025), YUME (Mao et al., 2025), and Astra—across three key dimensions: domain, control modality, and interaction horizon. In terms of domain, Astra is designed for general-purpose scenarios, showcasing their versatility in a wide range of applications. MatrixGame, however, is tailored specifically for game-related contexts, limiting its use to game-specific environments. YUME focuses on walking-specific domains, which is a niche area compared to the general-purpose methods. For control modalities, Wan-2.1 relies solely on text input, reflecting a language-driven paradigm. MatrixGame and YUME adopt traditional keyboard and mouse controls. Astra, enabled by MoAE, supports multiple modalities—including camera input, keyboard/mouse, and robot pose—providing more flexible and intuitive interaction. Finally, regarding the interaction horizon, Wan-2.1 and MatrixGame are limited to short spans of a few seconds. YUME extends this to 8–10 seconds, and Astra matches this longer horizon through noisy memory and the input-packing technique of (Zhang and Agrawala, 2025). Combined with its multi-modal control and general-purpose domain, Astra emerges as the most comprehensive solution among the compared methods.
@@
NVM (Bar et al., 2025) CDiT-XL $\sim$ 1B Full tuning
YUME (Mao et al., 2025) Wan2.1-14B $\sim$ 14B Full tuning
MatrixGame (He et al., 2025) Wan2.1-1.3B $\sim$ 1.8B Full tuning, cross-attn adapters
@@
Wan-2.1 (Wan et al., 2025) General Text A few seconds
MatrixGame (He et al., 2025) Game-specific Keyboard / Mouse A few seconds
YUME (Mao et al., 2025) Walking-specific Keyboard / Mouse 8-10 seconds
@@ C.2 Out-of-domain generalization
Figure B: Qualitative comparison with NVM (Bar et al., 2025). Astra consistently produces visually coherent results while faithfully following action inputs.
@@ C.3 Comparisons with more methods
To more comprehensively situate Astra within the broader landscape of visual world modeling, we provide extended quantitative and qualitative comparisons with additional state-of-the-art approaches. Specifically, we compare against methods such as Navigation World Models (NVM) (Bar et al., 2025) in Figure B and Table A, which are representative of recent approaches integrating multimodal inputs and action-conditioned predictions. Our results demonstrate that Astra achieves competitive or superior performance across key metrics, including visual fidelity, temporal consistency, and action responsiveness. While direct numerical comparisons are limited by differences in evaluation protocols and datasets, qualitative visualizations show that Astra generates long-horizon, coherent videos that faithfully follow user-specified actions, outperforming prior methods in capturing the causal dynamics of the environment. These comparisons highlight Astra’s ability to combine accurate action conditioning with long-term rollout stability, confirming its effectiveness as a general interactive world model across diverse scenarios. It is worth noting that many of the most advanced world models, such as Genie-3 (Bruce et al., 2024), are currently closed-source or provide only limited evaluation interfaces, making direct numerical comparison infeasible.
@@ C.4 Quantitative comparisons on larger datasets
To further address concerns about evaluation scale, we conduct an expanded study on a larger held-out set from the CityWalker (Liu et al., 2025) dataset. CityWalker is an egocentric urban navigation corpus collected from in-the-wild YouTube walking videos with pose estimates. We sample 100 scene images, each paired with its future ground-truth action trajectory, producing 100 long-horizon rollouts that robustly test cross-scene generalization and action adherence. Across this larger evaluation, Astra consistently achieves the best action-following accuracy and video quality, outperforming all baselines in metrics such as motion consistency, temporal stability, and VBench perceptual scores (Table D). These results confirm that Astra’s strong performance is not an artifact of small-set evaluation, and that the model maintains robust generalization and reliable action responsiveness even under substantially expanded test conditions.
@@
Method Instruction Following $\uparrow$ Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Wan-2.1 (Wan et al., 2025) 0.084 0.827 0.843 0.913 0.417 0.632 MatrixGame (He et al., 2025) 0.247 0.923 0.939 0.946 0.426 0.653 Yume (Mao et al., 2025) 0.619 0.933 0.927 0.972 0.511 0.628 Astra (Ours) 0.641 0.948 0.944 0.983 0.554 0.695
@@ Appendix D Discussion on visual inertia
Previous work (Zhang and Agrawala, 2025) has shown that longer video context can improve generation quality. However, in interactive world modeling, we observe that increasing context length can reduce action responsiveness. We term this phenomenon visual inertia—the model’s tendency to over-rely on past visual frames while neglecting user actions. To examine this, we train Astra with different history lengths without our proposed noisy-memory mechanism. As shown in Figure C, the action-following score drops substantially as context length increases, indicating that overly clean and long visual histories can dominate the model’s decision process. This motivates our noise-augmented memory design, which intentionally reduces visual dominance and encourages the model to integrate both historical context and action signals when generating future frames.